A particle moves along ay-plane with a position ofx (t) = sin(2t) + 3 cos(4t)y (t) = sin(t) – 6 cos(3t)a. Determine the position vector of the particle.b. Determine the velocity vector of the particle.c. Determine the acceleration vector of the particle.d. Determine the position vector of the particle when t = 4s.e. Determine the average acceleration from t = 0s to t= 5s.f. Determine the instantaneous acceleration when të Is.g. Determine the average velocity from t = ls to t= 3s.​

1
Expert's answer
2021-02-22T16:03:34-0500

a)


"\vec{r}=\begin{pmatrix}
 \sin(2t) + 3 \cos(4t) \\
 \sin(t) – 6 \cos(3t)
\end{pmatrix}"

b)


"\vec{v}=\begin{pmatrix}
 2 \cos(2t) -12\sin(4t) \\
 \cos(t) +18 \sin(3t)
\end{pmatrix}"

c)


"\vec{a}=\begin{pmatrix}
 -4 \sin(2t) -48\cos(4t) \\
 - \sin(t) +54 \cos(3t)
\end{pmatrix}"

d)


"\vec{r}=\begin{pmatrix}
 -1.88 \\
 - 5.82
\end{pmatrix}"

e)


"\vec{a}_{av}=\begin{pmatrix}
-2.93 \\
 2.20 
\end{pmatrix}"

f)


"\vec{a}(5)=\begin{pmatrix}
-17.4\\
-40.1 
\end{pmatrix}"

g)


"\vec{v}_{av}=\begin{pmatrix}
 1.65 \\
 -0.59
\end{pmatrix}"